OpenStudy (anonymous):

need help solving this equation step by step soybean meal is 18% protein, cornmeal is 9% protein how many lbs of each should be mixed together to get 360lb mixture that is 15%

OpenStudy (anonymous):

"soybean meal is 18% protein" does that mean soybean meal contains 18% protein?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Lets assume that we mixed x lbs of soybean and y lbs of cornmean

OpenStudy (anonymous):

So x + y = 360 lbs

OpenStudy (anonymous):

This is your first equation

OpenStudy (anonymous):

18/100(x)+9/100(y)=15/100

OpenStudy (anonymous):

18x+9y=15

OpenStudy (anonymous):

This is your second equation

OpenStudy (anonymous):

my mathlab says the second equation is 0.09x+0.18y=54

OpenStudy (anonymous):

this is where i started getting confused cant figure out where they got 54

OpenStudy (anonymous):

Excellent

OpenStudy (anonymous):

Matlab is right

OpenStudy (anonymous):

I missed something

OpenStudy (anonymous):

how did they get 54

OpenStudy (anonymous):

You see on the right hand side, I did it wrong

OpenStudy (anonymous):

18/100(x)+9/100(y)=(15/100)x(360)

OpenStudy (anonymous):

yeah thats y i asked about mathlab

OpenStudy (anonymous):

18/100(x)+9/100(y)=(15/100)x(360) =>0.18(x)+0.09(y)=54

OpenStudy (anonymous):

Got it?

OpenStudy (anonymous):

360(.15)=54

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

why any doubt?

OpenStudy (anonymous):

wasnt sure mathlab gives answer not explanation gets confusing to me im really really bad at math

OpenStudy (anonymous):

If you don't mind, is this problem for your kids

OpenStudy (anonymous):

I mean is this from their assignment

OpenStudy (anonymous):

no its for me, i go to strayer

OpenStudy (anonymous):

Thats very good.

OpenStudy (anonymous):

Any more problems at this time, which I can assist you with?

OpenStudy (anonymous):

yes complete the ordered triple 8x-y+z=32 3x+2y-3z=36 x-3y+2z=-14